OpenStudy (babyx3boo):

Solve for the following formula for X. Y=2mx+9b

OpenStudy (johnweldon1993):

\[\large y = 2mx + 9b\] Well, lets subtract 9b from both sides of the equation...we want to get 'x' by itself on one side So \[\large y - 9b = 2mx\] What would we do next?

OpenStudy (babyx3boo):

divide by 2m?

OpenStudy (johnweldon1993):

Perfect...and that would give us \[\large \frac{y - 9b}{2m} = x\] or just written correctly \[\large x = \frac{y - 9b}{2m}\]

OpenStudy (babyx3boo):

Thank you very much :)

OpenStudy (johnweldon1993):

No problem :)